# On Malle’s conjecture for the product of symmetric and nilpotent groups

### v2

@@ 3.2. Uniformity estimate for nilpotent groups
$\widetilde{\mathfrak{a}}_{i}\mathrel{\mathop{:}}=\prod_{j=1}^{r}\mathfrak{p}_{i_{j}}^{\lceil a_{i_{j}}/e_{i}\rceil}.$
@@
In this section, $G$ denotes a non-trivial finite nilpotent group. We also suppose that $G$ satisfies the appropriate conditions of Theorem 1.3 for $S_{3},S_{4},$ or $S_{5}$. We set
$\begin{split}\mathcal{F}_{k}(S_{n}\times G;X)\mathrel{\mathop{:}}=\{&(K,L)\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(K/k)\simeq S_{n},\\ &\operatorname{Gal}(L/k)\simeq G\text{ and }\mathrm{Disc}(KL)\leq X\}\end{split}$
and prove asymptotics for #$\#\mathcal{F}_{k}(S_{n}\times G;X)$ as $X\rightarrow\infty$. Then, it is easy to see that the same asymptotics holds for $N_{k}(S_{n}\times G;X)$. We first establish a lower bound.
@@
$N_{k}(S_{n}\times G;X)\geq$$\#\{(K,L)\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n}\text{ and }$
@@
The result then follows from the above bounds and Corollary 2.8 setting $\operatorname{inv}_{1}(K)\mathrel{\mathop{:}}=\mathrm{Disc}(K)$ and $\operatorname{inv}_{1}(L)\mathrel{\mathop{:}}=\mathrm{Disc}(L)$. ∎
@@
$\mathrm{Disc}_{Y}(KL)\mathrel{\mathop{:}}=\prod_{\mathfrak{p}\in\Omega_{k}}\mathrm{Disc}_{Y,\mathfrak{p}}(KL),$
@@
$\mathrm{Disc}_{Y,\mathfrak{p}}(KL)\mathrel{\mathop{:}}=\begin{cases}\mathrm{Disc}_{\mathfrak{p}}(KL)\text{ }\mathrm{if}\text{ }|\mathfrak{p}|\leq Y,\\ \mathrm{Disc}_{\mathfrak{p}}(K)^{|G|}\mathrm{Disc}_{\mathfrak{p}}(L)^{n}\text{}\mathrm{if}\text{ }|\mathfrak{p}|>Y.\end{cases}$
@@
(4.1) $\begin{split}N_{k,Y}(S_{n}\times G;X)\mathrel{\mathop{:}}=&\#\{(K,L)\mid[K\mathrel{\mathop{:}}k]=n,[L\mathrel{\mathop{:}}k]=|G|,\\ &\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\operatorname{Gal}(L/k)\simeq G\text{ and }\mathrm{Disc}_{Y}(KL)<X\}.\end{split}$
@@
Let $S_{Y}$ be the finite set of all primes of $k$ with absolute norm $|\mathfrak{p}|\leq Y$. Let $\Sigma_{Y}$ denote the set of tuples $(\mathcal{K}_{\mathfrak{p}},\mathcal{L}_{\mathfrak{p}})_{\mathfrak{p}\in S_{Y}}$ of pairs of local étale extensions over $k_{\mathfrak{p}}$ of degree $n$ and $|G|$ respectively. We will think of these tuples as specifying local conditions at all $\mathfrak{p}\in S_{Y}$ for $S_{n}$ and $G$-extensions. Accordingly, given a pair $(K,L)\in\mathcal{F}_{k}(S_{n}\times G;X)$ and a tuple $(\mathcal{K}_{\mathfrak{p}},\mathcal{L}_{\mathfrak{p}})_{\mathfrak{p}\in S_{Y}}\in\Sigma_{Y}$ we write $(K,L)\in((\mathcal{K}_{\mathfrak{p}},\mathcal{L}_{\mathfrak{p}}))_{\mathfrak{p}\in S_{Y}}$, if for each $\mathfrak{p}\in S_{Y}$,
@@
Next, we have the following Lemma relating $\mathrm{Disc}_{Y}$ and $\mathrm{Disc}$ under given local specifications,
@@
Let $Y>0$ and $\Pi\mathrel{\mathop{:}}=((\mathcal{K}_{\mathfrak{p}}),\mathcal{L}_{\mathfrak{p}}))_{\mathfrak{p}\in S_{Y}}\in\Sigma_{Y}$. If $(K,L)\in\Pi$ then,
$\mathrm{Disc}_{Y}(KL)\mathrel{\mathop{:}}=\frac{\mathrm{Disc}(K)^{|G|}\mathrm{Disc}(L)^{n}}{d_{\Pi}}$
@@
From the definition of $\mathrm{Disc}_{Y}$, it follows that,
$\mathrm{Disc}_{Y}(KL)\mathrel{\mathop{:}}=\prod_{\mathfrak{p}\in S_{Y}}\mathrm{Disc}_{\mathfrak{p}}(KL)\prod_{\mathfrak{p}\notin S_{Y}}\mathrm{Disc}_{\mathfrak{p}}(K)^{|G|}\mathrm{Disc}_{\mathfrak{p}}(L)^{n}.$
@@
$\mathrm{Disc}_{Y}(KL)\mathrel{\mathop{:}}=\frac{\mathrm{Disc}(K)^{|G|}\mathrm{Disc}(L)^{n}}{d_{\Pi}}$
@@
$N_{k,Y,\Pi}(S_{n}\times G;X)=$$\#\{(K,L)\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\operatorname{Gal}(L/k)\simeq G,$
@@
$N_{k,Y,\Pi}(S_{n}\times G;X)=$$\#\{(K,L)\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\operatorname{Gal}(L/k)\simeq G,$
@@
Given any $G$-extension $L\in\Pi$, we also consider the counting function,
$N_{k,Y,\Pi}^{L}(S_{n}\times G;X)=$$\#\{K\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},(K,L)\in\Pi,$
@@
$\sum_{\begin{subarray}{c}L\in\Pi,\\ \mathrm{Disc}(L)>Z\end{subarray}}N_{k,Y,\Pi}^{L}(S_{n}\times G;X)\ll_{k,S_{n}\times G,\Pi}\left(\sum_{\begin{subarray}{c}L\in\Pi,\\ \mathrm{Disc}(L)>Z\end{subarray}}\frac{1}{\mathrm{Disc}(L)^{n/|G|}}\right)X^{\frac{1}{|G|}}.$
@@
$S_{\sigma}\mathrel{\mathop{:}}=\{s\geq 0\mathrel{\mathop{:}}N_{k,\mathfrak{q},\sigma}(S_{n};X)\ll_{\epsilon}X/|\mathfrak{q}|^{s-\epsilon}\text{ for all squarefree }\mathfrak{q}\text{ in }k\},$
@@
Using computations in Proposition 2.4 and bounds on $s_{\sigma}$ from Theorem 3.1 we conclude that the above inequality holds. ∎
Remark 4.7.
We use the fact that the constants $s_{\sigma}$ are not smaller than the corresponding constants that appear in Theorem 3.1 for suitable $\sigma$, we denote the corresponding constants by $r_{\sigma}$. For other conjugacy classes $\sigma$ not covered in Theorem 3.1 we use the fact that the constant $s_{\sigma}$ is non-negative, in this case, we define $r_{\sigma}$ to be zero.
@@ 4.0.1. The difference $N_{k}(S_{n}\times G;X)-N_{k,Y}(S_{n}\times G;X)$
(4.3) $\begin{split}&N_{k}(S_{n}\times G;X)-N_{k,Y}(S_{n}\times G;X)\\ =&\#\{(K,L)\mid[K\mathrel{\mathop{:}}k]=n,[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\\ &\operatorname{Gal}(L/k)\simeq G,\text{ and }\mathrm{Disc}(KL)<X<\mathrm{Disc}_{Y}(KL)\}.\end{split}$
@@
Let $\Sigma_{k}$ denote the set of all standard tuples for the number field $k$. Given $(K,L)$ and a specification of local conditions $\Pi\mathrel{\mathop{:}}=(S,\mathcal{L})\in\Sigma_{k}$ as above, we write $(K,L)\in\Pi$ if following conditions are satisfied.
@@
Let $\Pi\mathrel{\mathop{:}}=(S,\mathcal{L})\in\Sigma_{k}$ be a standard tuple and suppose that $(K,L)\in\Pi$ then,
$\mathrm{Disc}(KL)\mathrel{\mathop{:}}=\frac{\mathrm{Disc}(K)^{|G|}\mathrm{Disc}(L)^{n}}{d_{\Pi}}$
@@
Let $\mathfrak{p}\notin S$ then either $K/k$ or $L/k$ is not ramified at $\mathfrak{p}$. Without loss of generality, we assume that $K/k$ is not ramified at $\mathfrak{p}$. We deduce from the upper bound in Proposition 2.1 that the $\mathrm{disc}_{\mathfrak{p}}(KL/k)\mid\mathrm{disc}_{\mathfrak{p}}(L/k)^{n}$. We recall that in the tower $KL/L/k$ we have the following formula,
@@
(4.4) $N_{k,\Pi}(S_{n}\times G;X)\mathrel{\mathop{:}}=\#\{(K,L)\in\mathcal{E}_{k}\mathrel{\mathop{:}}(K,L)\in\Pi\text{ and }\mathrm{Disc}(KL)<X\},$
@@
(4.5) $\mathcal{E}_{k}\mathrel{\mathop{:}}=\{(K,L)\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n}\text{ and }\operatorname{Gal}(L/k)\simeq G\}.$
@@
(4.6) $\begin{split}N_{k,\Pi}(S_{n}\times G;X)=&\#\{(K,L)\in\mathcal{E}_{k}\mathrel{\mathop{:}}(K,L)\in\Pi\text{ and }\\ &\mathrm{Disc}(K)^{|G|}\mathrm{Disc}(L)^{n}<Xd_{\Pi}\}\\ =&\#\{(K,L)\in\mathcal{E}_{k}\mathrel{\mathop{:}}(K,L)\in\Pi\text{ and }\\ &\prod_{\mathfrak{p}\notin S}\mathrm{Disc}_{\mathfrak{p}}(K)^{|G|}\mathrm{Disc}_{\mathfrak{p}}(L)^{n}<{X}/{\prod_{\mathfrak{p}\in S}|\mathfrak{p}|^{\exp(h_{\mathfrak{p}},g_{\mathfrak{p}})}}\}\end{split}$
@@
$\mathrm{Disc}_{res}^{\Pi}(K)\mathrel{\mathop{:}}=\prod_{\mathfrak{p}\notin S}\mathrm{Disc}_{\mathfrak{p}}(K).$
@@
$N_{k,\Pi,res}(S_{n};X)\mathrel{\mathop{:}}=\#\{K\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,\operatorname{Gal}(K/k)\simeq S_{n},K\in\Pi,\text{ and }\mathrm{Disc}_{res}^{\Pi}(K)\leq X\},$
@@
$N_{k,\Pi,res}(G;X)\mathrel{\mathop{:}}=\#\{L\mathrel{\mathop{:}}[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(L/k)\simeq G,L\in\Pi,\text{ and }\mathrm{Disc}_{res}^{\Pi}(L)\leq X\}.$
Set $C(S_{n})$ to be the set of conjugacy classes in $S_{n}$. Given $\sigma\in C(S_{n})$, we take $\mathfrak{q}_{\sigma}\mathrel{\mathop{:}}=\prod_{\mathfrak{p}\in S,h_{\mathfrak{p}}=\sigma}\mathfrak{p}$. Recall that $r_{\sigma}$ is defined in Lemma 4.6.
@@
Using definition of $N_{k,\Pi,res}(S_{n};X)$ we note that,
(4.9) $\begin{split}N_{k,\Pi,res}(S_{n};X)&\leq\#\{K\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,\operatorname{Gal}(K/k)\simeq S_{n},K\in\Pi,\text{ and }\\ &\mathrm{Disc}(K)\leq X\prod_{\mathfrak{p}\in S}|\mathfrak{p}|^{\exp(h_{\mathfrak{p}})}\}\\ &\leq\#\{K\mathrel{\mathop{:}}[K\mathrel{\mathop{:}}k]=n,\operatorname{Gal}(K/k)\simeq S_{n},K\in\Pi,\text{ and }\\ &\mathrm{Disc}(K)\leq CX\prod_{\sigma\in C(S_{n})}|\mathfrak{q}_{\sigma}|^{\mathrm{ind}(\sigma)}\},\end{split}$
@@
The proof is a simple application of Proposition 3.6. Note that the following inequality holds,
$N_{k,\Pi,res}(G;X)\leq\#\{L\mathrel{\mathop{:}}[L\mathrel{\mathop{:}}k]=|G|,\operatorname{Gal}(L/k)\simeq G,\text{ and }\mathrm{Disc}(L)\leq X\prod_{\mathfrak{p}\in S}|\mathfrak{p}|^{\exp(g_{\mathfrak{p}})}\}.$
@@
$\delta\mathrel{\mathop{:}}=\max_{\sigma\in C(S_{n}),g\in G}(-r_{\sigma}+\mathrm{ind}(\sigma)-\mathrm{ind}(\sigma,g)/|G|).$
@@
$N_{k,\Pi}(S_{n}\times G;X)=$$\#\{(K,L)\in\mathcal{E}_{k}\mathrel{\mathop{:}}(K,L)\in\Pi\text{ and }$
@@
$=$ #$\#\{(K,L)\in\mathcal{E}_{k}\mathrel{\mathop{:}}(K,L)\in\Pi\text{ and }$
@@
We have previously established upper bounds on the number of suitable $S_{n}$ and $G$-extensions when ordered by the invariant $\mathrm{Disc}_{res}$. Therefore, a straightforward application of Corollary 2.8 with $\mathcal{P}_{1},\mathcal{P}_{2}$ properties that $K\in\Pi$, $L\in\Pi$ respectively and $\operatorname{inv}_{1},\operatorname{inv}_{2}$ are restricted discriminant together with Lemmas 4.13 and 4.14 yields the desired bound on $N_{k,\Pi}(S_{n}\times G;X)$. ∎

### v3

@@ 3.2. Uniformity estimate for nilpotent groups
$\widetilde{\mathfrak{a}}_{i}:=\prod_{j=1}^{r}\mathfrak{p}_{i_{j}}^{\lceil a_{i_{j}}/e_{i}\rceil}.$
@@
In this section, $G$ denotes a non-trivial finite nilpotent group. The results in this section are analogous to those in [Wan21, Section 5]. We also suppose that $G$ satisfies the appropriate conditions of Theorem 1.3 for $S_{3},S_{4},$ or $S_{5}$. We set
$\begin{split}\mathcal{F}_{k}(S_{n}\times G;X):=\{&(K,L):[K:k]=n,[L:k]=|G|,\operatorname{Gal}(K/k)\simeq S_{n},\\ &\operatorname{Gal}(L/k)\simeq G\text{ and }\mathrm{Disc}(KL)\leq X\}\end{split}$
and prove asymptotics for #$\#\mathcal{F}_{k}(S_{n}\times G;X)$ as $X\rightarrow\infty$. We shall prove that the same asymptotics hold for $N_{k}(S_{n}\times G;X)$. We first establish a lower bound.
@@
$N_{k}(S_{n}\times G;X)\geq$$\#\{(K,L):[K:k]=n,[L:k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n}\text{ and }$
@@
The result then follows from the above bounds and Corollary 2.8 setting $\operatorname{inv}_{1}(K):=\mathrm{Disc}(K)$ and $\operatorname{inv}_{1}(L):=\mathrm{Disc}(L)$. ∎
@@
$\mathrm{Disc}_{Y}(KL):=\prod_{\mathfrak{p}\in\Omega_{k}}\mathrm{Disc}_{Y,\mathfrak{p}}(KL),$
@@
$\mathrm{Disc}_{Y,\mathfrak{p}}(KL):=\begin{cases}\mathrm{Disc}_{\mathfrak{p}}(KL)\text{ }\mathrm{if}\text{ }|\mathfrak{p}|\leq Y,\\ \mathrm{Disc}_{\mathfrak{p}}(K)^{|G|}\mathrm{Disc}_{\mathfrak{p}}(L)^{n}\text{}\mathrm{if}\text{ }|\mathfrak{p}|>Y.\end{cases}$
@@
(4.1) $\begin{split}N_{k,Y}(S_{n}\times G;X):=&\#\{(K,L)\mid[K:k]=n,[L:k]=|G|,\\ &\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\operatorname{Gal}(L/k)\simeq G\text{ and }\mathrm{Disc}_{Y}(KL)<X\}.\end{split}$
Remark 4.3.
We note here that a similar definition arises in Wang’s work. The above definition generalizes [Wan21, equation (5.4)].
@@
Let $S_{Y}$ be the finite set of all primes of $k$ with absolute norm $|\mathfrak{p}|\leq Y$. Let $\Sigma_{Y}$ denote the set of tuples $(\mathcal{K}_{\mathfrak{p}},\mathcal{L}_{\mathfrak{p}})_{\mathfrak{p}\in S_{Y}}$ of pairs of local étale extensions over $k_{\mathfrak{p}}$ of degree $n$ and $|G|$ respectively. We will think of these tuples as specifying local conditions at all $\mathfrak{p}\in S_{Y}$ for $S_{n}$ and $G$-extensions. This argument is indeed an analogue of [Wan21, p.116, l. -12]. Accordingly, given a pair $(K,L)\in\mathcal{F}_{k}(S_{n}\times G;X)$ and a tuple $(\mathcal{K}_{\mathfrak{p}},\mathcal{L}_{\mathfrak{p}})_{\mathfrak{p}\in S_{Y}}\in\Sigma_{Y}$ we write $(K,L)\in((\mathcal{K}_{\mathfrak{p}},\mathcal{L}_{\mathfrak{p}}))_{\mathfrak{p}\in S_{Y}}$, if for each $\mathfrak{p}\in S_{Y}$,
@@
Next, we have the following Lemma relating $\mathrm{Disc}_{Y}$ and $\mathrm{Disc}$ under given local specifications.
@@
Let $Y>0$ and $\Pi:=((\mathcal{K}_{\mathfrak{p}}),\mathcal{L}_{\mathfrak{p}}))_{\mathfrak{p}\in S_{Y}}\in\Sigma_{Y}$. If $(K,L)\in\Pi$ then,
$\mathrm{Disc}_{Y}(KL):=\frac{\mathrm{Disc}(K)^{|G|}\mathrm{Disc}(L)^{n}}{d_{\Pi}}$
@@
The results is similar to Wang’s argument (cf. [Wan21, equation (5.5)]), which applies in the case when $G$ is abelian. From the definition of $\mathrm{Disc}_{Y}$, it follows that,
$\mathrm{Disc}_{Y}(KL):=\prod_{\mathfrak{p}\in S_{Y}}\mathrm{Disc}_{\mathfrak{p}}(KL)\prod_{\mathfrak{p}\notin S_{Y}}\mathrm{Disc}_{\mathfrak{p}}(K)^{|G|}\mathrm{Disc}_{\mathfrak{p}}(L)^{n}.$
@@
$\mathrm{Disc}_{Y}(KL):=\frac{\mathrm{Disc}(K)^{|G|}\mathrm{Disc}(L)^{n}}{d_{\Pi}}$
@@
$N_{k,Y,\Pi}(S_{n}\times G;X)=$$\#\{(K,L):[K:k]=n,[L:k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\operatorname{Gal}(L/k)\simeq G,$
@@
$N_{k,Y,\Pi}(S_{n}\times G;X)=$$\#\{(K,L):[K:k]=n,[L:k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\operatorname{Gal}(L/k)\simeq G,$
@@
[…]
$N_{k,Y,\Pi}^{L}(S_{n}\times G;X)=$$\#\{K:[K:k]=n,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},(K,L)\in\Pi,$
@@
$\sum_{\begin{subarray}{c}L\in\Pi,\\ \mathrm{Disc}(L)>Z\end{subarray}}N_{k,Y,\Pi}^{L}(S_{n}\times G;X)\ll_{k,S_{n}\times G,\Pi,Y}\left(\sum_{\begin{subarray}{c}L\in\Pi,\\ \mathrm{Disc}(L)>Z\end{subarray}}\frac{1}{\mathrm{Disc}(L)^{n/|G|}}\right)X^{\frac{1}{|G|}}.$
@@
$S_{\sigma}:=\{s\geq 0:N_{k,\mathfrak{q},\sigma}(S_{n};X)\ll_{\epsilon}X/|\mathfrak{q}|^{s-\epsilon}\text{ for all squarefree }\mathfrak{q}\text{ in }k\},$
@@
In the case when $\sigma$ is totally ramified (or overramified), the result follows from Theorem 3.1. In all other cases, the result is a direct consequence of Proposition 2.5 by simply employing the trivial lower bound $s_{\sigma}\geq 0$. Note that when $n=5$, and $\sigma$ is conjugate to $(12345)$, then one happens to be in the totally ramified case and Theorem 3.1 applies. ∎
@@ 4.0.1. The difference $N_{k}(S_{n}\times G;X)-N_{k,Y}(S_{n}\times G;X)$
(4.3) $\begin{split}&N_{k}(S_{n}\times G;X)-N_{k,Y}(S_{n}\times G;X)\\ =&\#\{(K,L)\mid[K:k]=n,[L:k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n},\\ &\operatorname{Gal}(L/k)\simeq G,\text{ and }\mathrm{Disc}(KL)<X<\mathrm{Disc}_{Y}(KL)\}.\end{split}$
@@
Let $\Sigma_{k}$ denote the set of all standard tuples for the number field $k$. Given $(K,L)$ and a specification of local conditions $\Pi:=(S,\mathcal{L})\in\Sigma_{k}$ as above, we write $(K,L)\in\Pi$ if following conditions are satisfied.
@@
Remark 4.9.
We note that the notion of standard tuples arises in Wang’s work; $\Pi$ used in this context is the exact analogue of the set considered in [Wan21, equation (5.10)].
@@
Let $\Pi:=(S,\mathcal{L})\in\Sigma_{k}$ be a standard tuple and suppose that $(K,L)\in\Pi$ then,
$\mathrm{Disc}(KL):=\frac{\mathrm{Disc}(K)^{|G|}\mathrm{Disc}(L)^{n}}{d_{\Pi}}$
@@
This result generalizes [Wan21, equation (5.11)]. Let $\mathfrak{p}\notin S$ then either $K/k$ or $L/k$ is not ramified at $\mathfrak{p}$. Without loss of generality, we assume that $K/k$ is not ramified at $\mathfrak{p}$. We deduce from the upper bound in Proposition 2.1 that the $\mathrm{disc}_{\mathfrak{p}}(KL/k)\mid\mathrm{disc}_{\mathfrak{p}}(L/k)^{n}$. We recall that in the tower $KL/L/k$ we have the following formula,
@@
(4.4) $N_{k,\Pi}(S_{n}\times G;X):=\#\{(K,L)\in\mathcal{E}_{k}:(K,L)\in\Pi\text{ and }\mathrm{Disc}(KL)<X\},$
@@
(4.5) $\mathcal{E}_{k}:=\{(K,L):[K:k]=n,[L:k]=|G|,\operatorname{Gal}(\widetilde{K}/k)\simeq S_{n}\text{ and }\operatorname{Gal}(L/k)\simeq G\}.$
@@
(4.6) $\begin{split}N_{k,\Pi}(S_{n}\times G;X)=&\#\{(K,L)\in\mathcal{E}_{k}:(K,L)\in\Pi\text{ and }\\ &\mathrm{Disc}(K)^{|G|}\mathrm{Disc}(L)^{n}<Xd_{\Pi}\}\\ =&\#\{(K,L)\in\mathcal{E}_{k}:(K,L)\in\Pi\text{ and }\\ &\prod_{\mathfrak{p}\notin S}\mathrm{Disc}_{\mathfrak{p}}(K)^{|G|}\mathrm{Disc}_{\mathfrak{p}}(L)^{n}<{X}/{\prod_{\mathfrak{p}\in S}|\mathfrak{p}|^{\exp(h_{\mathfrak{p}},g_{\mathfrak{p}})}}\}\end{split}$
@@
$\mathrm{Disc}_{res}^{\Pi}(K):=\prod_{\mathfrak{p}\notin S}\mathrm{Disc}_{\mathfrak{p}}(K).$
@@
$N_{k,\Pi,res}(S_{n};X):=\#\{K:[K:k]=n,\operatorname{Gal}(K/k)\simeq S_{n},K\in\Pi,\text{ and }\mathrm{Disc}_{res}^{\Pi}(K)\leq X\},$
@@
$N_{k,\Pi,res}(G;X):=\#\{L:[L:k]=|G|,\operatorname{Gal}(L/k)\simeq G,L\in\Pi,\text{ and }\mathrm{Disc}_{res}^{\Pi}(L)\leq X\}.$
Set $C(S_{n})$ to be the set of conjugacy classes in $S_{n}$. Given $\sigma\in C(S_{n})$, we take $\mathfrak{q}_{\sigma}:=\prod_{\mathfrak{p}\in S,h_{\mathfrak{p}}=\sigma}\mathfrak{p}$. Recall that $r_{\sigma}$ is defined in Lemma 4.7.
@@
This result is adaptation of [Wan21, equation (5.13)]. Using definition of $N_{k,\Pi,res}(S_{n};X)$ we note that,
(4.9) $\begin{split}N_{k,\Pi,res}(S_{n};X)&\leq\#\{K:[K:k]=n,\operatorname{Gal}(K/k)\simeq S_{n},K\in\Pi,\text{ and }\\ &\mathrm{Disc}(K)\leq X\prod_{\mathfrak{p}\in S}|\mathfrak{p}|^{\exp(h_{\mathfrak{p}})}\}\\ &\leq\#\{K:[K:k]=n,\operatorname{Gal}(K/k)\simeq S_{n},K\in\Pi,\text{ and }\\ &\mathrm{Disc}(K)\leq CX\prod_{\sigma\in C(S_{n})}|\mathfrak{q}_{\sigma}|^{\mathrm{ind}(\sigma)}\},\end{split}$
@@
Our result is proven as a generalization of [Wan21, equation (5.14)]. The proof is a simple application of Proposition 3.6. Note that the following inequality holds,
$N_{k,\Pi,res}(G;X)\leq\#\{L:[L:k]=|G|,\operatorname{Gal}(L/k)\simeq G,\text{ and }\mathrm{Disc}(L)\leq X\prod_{\mathfrak{p}\in S}|\mathfrak{p}|^{\exp(g_{\mathfrak{p}})}\}.$
@@
$\delta:=\max_{\sigma\in C(S_{n}),g\in G}(-r_{\sigma}+\mathrm{ind}(\sigma)-\mathrm{ind}(\sigma,g)/|G|).$
@@
$N_{k,\Pi}(S_{n}\times G;X)=$$\#\{(K,L)\in\mathcal{E}_{k}:(K,L)\in\Pi\text{ and }$
@@
$=$ #$\#\{(K,L)\in\mathcal{E}_{k}:(K,L)\in\Pi\text{ and }$
@@
We have previously established upper bounds on the number of suitable $S_{n}$ and $G$-extensions when ordered by the invariant $\mathrm{Disc}_{res}$. Therefore, a straightforward application of Corollary 2.8 with $\mathcal{P}_{1},\mathcal{P}_{2}$ properties that $K\in\Pi$, $L\in\Pi$ respectively and $\operatorname{inv}_{1},\operatorname{inv}_{2}$ are restricted discriminant together with Lemmas 4.14 and 4.15 yields the desired bound on $N_{k,\Pi}(S_{n}\times G;X)$. ∎